OpenStudy (anonymous):

Given a circle with radius r = 4 and a intercepted arc length of 5π/3, what is the area of the related sector (in radians)?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

\[=\pi r ^{2} \frac{5\pi/3}{2 \pi}\]

OpenStudy (anonymous):

is that the final answer?

OpenStudy (dumbcow):

no you need to replace r with 4 and simplify

OpenStudy (anonymous):

so it would be 3.14(4)(4)?

OpenStudy (dumbcow):

yes then the other part would reduce to 5/6

OpenStudy (anonymous):

ok i got 50.24 5(pi)/6?

OpenStudy (anonymous):

is that correct

OpenStudy (dumbcow):

yes

OpenStudy (dumbcow):

wait no take away the pi

OpenStudy (anonymous):

do i multiply 50.24 by 5/6?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

ok got that but its not in radians